OpenStudy (aaronandyson):

a stone is dropped in to a well 20 deep.another stone is dropped 1 seconds later..if both reach the water surface simultaneously ,v is equal to?

OpenStudy (aaronandyson):

take g = 10 ms^-2

OpenStudy (aaronandyson):

Optiona: 30 15 20 10

OpenStudy (evoker):

Ah i assume you are saying the first is thrown with some initial velocity.

OpenStudy (aaronandyson):

I have typed the entire question.

OpenStudy (evoker):

dropped usually means they have the same initial velocity of 0 so unless you are figuring in air drag, The second should hit the water 1 second later than the first.

OpenStudy (aaronandyson):

nothing is mentioned...so u = 0?

OpenStudy (evoker):

But based on your question I would have to assume one has an initial velocity that isn't zero so that they can both hit at the same time.

OpenStudy (welshfella):

yea thats correct

OpenStudy (welshfella):

if initial velocity is zero for both stones they will not hit the water at the same time.

OpenStudy (evoker):

So in that case I would assume the second is perhaps dropped and the first is thrown up, or alternatively the first is dropped and the second is thrown down.

OpenStudy (welshfella):

if we apply the formula v^2 = u^2 + 2as v^2 = 0^2 + 2*10*20 v = sqrt400

OpenStudy (aaronandyson):

v = 20ms^-1 then...but thats incorrect...

OpenStudy (evoker):

Sure you could go that way, hes not done, he just saying that the stone dropped with an initial velocity of 0 will have a final speed of 20, and using vf=vi+at you could solve that it occurs in 2 seconds.

OpenStudy (welshfella):

well thats is correct for a stone dropped with zero initial velocity

OpenStudy (aaronandyson):

but here we have take initial v as u right?

OpenStudy (evoker):

So then you could solve for vi for a stone that drops 20 in 1 seconds or 3 seconds depending on the setup of the problem.

OpenStudy (evoker):

or u as you call it

OpenStudy (evoker):

lets see if you use t=1 you would have 20=v*1+1/2*10*1^2, which yields 20=v+5 so if they throw it down at 15m/s, the second would catch up

OpenStudy (welshfella):

the question as it stands does not make sense. The second stone must have u>0 for them to arrive at bottom at same time.

OpenStudy (evoker):

or alternatively if t=3 you would have 20=v*3+1/2*10*3^2. 20=3v+45 -25=3v or v=25/3 thrown up, but ultimately welshfella is correct, really need the original problem as this one doesn't make sense.

OpenStudy (evoker):

Though from your options I would lean towards the 15.

OpenStudy (aaronandyson):

lemme redo it s = ut + 0.5st^2 20 = u + 0.5*10*1 20 = u + 5 u = 15?

OpenStudy (evoker):

That would be if we are looking at the case where the first is dropped and the second is thrown down, and the second s in the first line would be an a.